OpenStudy (anonymous):

To reduce waste, a local farm offers two reusable gallon milk jugs. One costs twenty dollars, and you can get it filled with fresh milk for two dollars. The other costs fifteen dollars, and you can get it filled with fresh milk for two point five zero dollars Which graph represents the system of linear equations for this problem? To reduce waste, a local farm offers two reusable gallon milk jugs. One costs twenty dollars, and you can get it filled with fresh milk for two dollars. The other costs fifteen dollars, and you can get it filled with fresh milk for two point five zero dollars Which graph represents the system of linear equations for this problem? @Mathematics

OpenStudy (anonymous):

im going to attach the graphs.

OpenStudy (turingtest):

red: 15+2.5x blue:20+2x

OpenStudy (turingtest):

are you asking which graph is correct?

OpenStudy (anonymous):

yes pleas e im sorry i had to attach them all one by one and there is one left

OpenStudy (turingtest):

well it is the first one you posted that represents the cost blue:cost=20+2x red: cost=15+2.5x

OpenStudy (turingtest):

04_pp_01_036.jpg this one

OpenStudy (anonymous):

kk thanks alot